OpenStudy (anonymous):

In a right triangle, 3/5 of the length of the longer leg is equal to 1/2 of the length of the hypotenuse? What is the ratio of the length of the hypotenuse to the length of the shorter leg?

OpenStudy (faman39):

@mattfeury, can u help?

OpenStudy (anonymous):

who's that?

OpenStudy (faman39):

@myininaya, can u help with it?

OpenStudy (faman39):

they are gr8 in math :)

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

where r they

OpenStudy (faman39):

@igbasallote, can u help with it?

OpenStudy (faman39):

they are gr8 OS moderator and users

OpenStudy (faman39):

U dont know them?

OpenStudy (anonymous):

no

OpenStudy (faman39):

@vikayefremova

OpenStudy (anonymous):

can u tell them to come? they're not responding

OpenStudy (faman39):

No, they are busy rite now, but u will get soon :) dont worry

OpenStudy (anonymous):

sorry, i'm not so good in math :(

OpenStudy (anonymous):

hey sami-21, u know how to do it?

OpenStudy (anonymous):

if u do, type here. I'll brb. I will give u medals if u show me how to do it :)

OpenStudy (anonymous):

let length of hypotenuse is h let length of longer side is a let length of shorter side is b (3/5)h=(1/2)a so h=(5/6)a now ratio of h/b is =(5/6)a/b

OpenStudy (anonymous):

oops i made a mistake!

OpenStudy (anonymous):

let length of hypotenuse is h let length of longer side is a let length of shorter side is b (1/2)h=(3/5)a so h=(6/5)a now ratio of h/b is =(6/5)a/b

OpenStudy (anonymous):

so what's the answer?

OpenStudy (anonymous):

?